How much water (ounces)must be evaporated from 16 ounces of a 2% salt solution to make a 4% salt solution?

Answer by stanbon(75887)

How much water (ounces)must be evaporated from 16 ounces of a 2% salt solution to make a 4% salt solution?
---------------------
2% solution DATA:
Amt. = 16 ounces ; salt = 0.02*16 = 0.32 ounces
-------------------------
Water DATA:
Amt. = x ounces ; salt = 0
---------------------------
Resulting Mixture DATA
Amt. = 16-x ; salt = 0.04(16-x) = 0.64 - 0.04x ounces
--------------------
EQUATION:
salt - salt = salt
0.32 - 0 = 0.64 - 0.04x
0.32 = 0.04x
x = 8 ounces of water must be evaporated
=================
Cheers,
Stan H.